# A high-order procedure for computing globally optimal Wannier functions in one-dimensional crystalline insulators

### v1

@@
September 5, 2024
$\mbox{}^{\dagger}$ Department of Mathematics, UC Davis
$\mbox{}^{\ddagger}$ Department of Mathematics, Yale University
@@ 2.1 Notation
We let $\ell^{2}$ denote the Hilbert space for complex sequences $\mathbf{x}=\left\{x_{i}\right\}_{i=-\infty}^{\infty}$ such that $\sum_{i=-\infty}^{\infty}\absolutevalue{x}^{2}<\infty$ with the standard inner product. We denote the right shift operator in $\ell^{2}$ by $R$, where
@@ 2.5 Derivatives of eigenvalues and eigenvectors
$\frac{\mathrm{d}\lambda(z_{0})}{\mathrm{d}z}=\mathbf{v}^{*}(z_{0})\frac{\mathrm{d}T(z_{0})}{\mathrm{d}z}\mathbf{v}(z_{0}),$ (27)
@@
$\frac{\mathrm{d}\mathbf{v}(z_{0})}{\mathrm{d}z}=\sum_{\begin{subarray}{c}j\\ \lambda_{j}\neq\lambda\end{subarray}}\mathbf{v}_{j}(z_{0})\frac{\mathbf{v}^{*}_{j}(z_{0})\frac{\mathrm{d}T(z_{0})}{\mathrm{d}z}\mathbf{v}(z_{0})}{\lambda(z_{0})-\lambda_{j}(z_{0})},$ (30)
@@ 3.3 Perturbation analysis in the Fourier domain
$\frac{\mathrm{d}}{\mathrm{d}k}E(k)=\mathbf{y}^{*}(k)\mathbf{S}(k)\mathbf{y}(k),$ (60)
@@
$\mathbf{y}^{*}(k)\frac{\mathrm{d}}{\mathrm{d}k}\mathbf{y}(k)=0,$ (64)
@@ 4 Constructing optimally localized Wannier functions
There are various schemes resembling the one above. For example, [20] and Section IV.C in [15] describe a conceptually similar algorithm. Instead of solving the ODE in the second stage above, $u_{k}$, for varying values of $k$ on a grid in $I_{\mathrm{bz}}$, are computed independently and an approximate SVD-based approach is used to fix the broken phase. For a large number of grid points, the SVD-based approach should yield a decent approximation to the ODE solution above (see Remark 4.8 for further discussion). However, the authors in [15, 20] did not prove the solution’s optimality as stated in Corollary 3.14. Another similar approach can be found in Section III in [4], where an approximate parallel transport scheme was used to construct $u_{k}$. Different matching conditions are imposed to obtain assignments of $u_{k}$ that are only continuous in $k$, producing Wannier functions that decay as $1/x^{2}$, which is then used as inputs for the general local optimization approach in [15].
@@ 4.1.3 Solving the initial value problem
$E^{\prime}(k)$ $=\mathbf{y}(k)^{*}\mathbf{S}(k)\mathbf{y}(k)$ (115)
@@ 6.1 Gaussian potential
In this experiment, we set $a=1$ (so $\Omega=2\pi$), $M=10$, and $\ell=1,2,3$. The potential is taken to be

### v2

@@
September 12, 2024
$\mbox{}^{\dagger}$ Department of Mathematics, UC Davis, Davis, CA 95616
$\mbox{}^{\ddagger}$ Department of Mathematics, Yale University, New Haven, CT 06511
@@ 2.1 Notation
We let $\ell^{2}$ denote the Hilbert space for complex sequences $\mathbf{x}=\left\{x_{i}\right\}_{i=-\infty}^{\infty}$ such that $\sum_{i=-\infty}^{\infty}\absolutevalue{x_{i}}^{2}<\infty$ with the standard inner product. We denote the right shift operator in $\ell^{2}$ by $R$, where
@@ 2.5 Derivatives of eigenvalues and eigenvectors
$\frac{\mathrm{d}\lambda(z_{0})}{\mathrm{d}z}=\mathbf{v}(z_{0})^{*}\frac{\mathrm{d}T(z_{0})}{\mathrm{d}z}\mathbf{v}(z_{0}),$ (27)
@@
$\frac{\mathrm{d}\mathbf{v}(z_{0})}{\mathrm{d}z}=\sum_{\begin{subarray}{c}j\\ \lambda_{j}\neq\lambda\end{subarray}}\mathbf{v}_{j}(z_{0})\frac{\mathbf{v}_{j}(z_{0})^{*}\frac{\mathrm{d}T(z_{0})}{\mathrm{d}z}\mathbf{v}(z_{0})}{\lambda(z_{0})-\lambda_{j}(z_{0})},$ (30)
@@ 3.3 Perturbation analysis in the Fourier domain
$\frac{\mathrm{d}}{\mathrm{d}k}E(k)=\frac{\mathbf{y}(k)^{*}\mathbf{S}(k)\mathbf{y}(k)}{\mathbf{y}(k)^{*}\mathbf{y}(k)},$ (60)
@@
$\mathbf{y}(k)^{*}\frac{\mathrm{d}}{\mathrm{d}k}\mathbf{y}(k)=0,$ (64)
@@ 4 Constructing optimally localized Wannier functions
There are various schemes resembling the one above. For example, [20] and Section IV.C in [15] describe a conceptually similar algorithm. Instead of solving the ODE in the second stage above, $u_{k}$, for varying values of $k$ on a grid in $I_{\mathrm{bz}}$, are computed independently and an approximate SVD-based approach is used to fix the broken phase. For a large number of grid points, the SVD-based approach should yield a decent approximation to the ODE solution above (see Remark 4.8 for further discussion). However, the authors in [15, 20] did not prove the solution’s optimality as stated in Corollary 3.14. Another similar approach can be found in Section III in [4], where an approximate parallel transport scheme was used to construct $u_{k}$. Different matching conditions are imposed to obtain assignments of $u_{k}$ that are only continuous in $k$, producing Wannier functions that decay as $1/x^{2}$, which are then used as inputs for the general local optimization approach in [15].
@@ 4.1.3 Solving the initial value problem
$E^{\prime}(k)$ $=\frac{\mathbf{y}(k)^{*}\mathbf{S}(k)\mathbf{y}(k)}{\mathbf{y}(k)^{*}\mathbf{y}(k)}$ (115)
@@ 6.1 Gaussian potential
In this experiment, we set $a=2\pi$ (so $\Omega=1$), $M=10$, and $\ell=1,2,3$. The potential is taken to be
@@ 7 Conclusions
Acknowledgements
The authors thank John Schotland for drawing their attention to this subject. The authors also thank Vladimir Rokhlin for helpful discussions pertaining to this work.